# Latent learning: episodic memory complements parametric learning by enabling flexible reuse of experiences

### v1

@@ Abstract
When do machine learning systems fail to generalize, and what mechanisms could improve their generalization? Here, we draw inspiration from cognitive science to argue that one weakness of machine learning systems is their failure to exhibit latent learning—learning information that is not relevant to the task at hand, but that might be useful in a future task. We show how this perspective links failures ranging from the reversal curse in language modeling to new findings on agent-based navigation. We then highlight how cognitive science points to episodic memory as a potential part of the solution to these issues. Correspondingly, we show that a system with an oracle retrieval mechanism can use learning experiences more flexibly to generalize better across many of these challenges. We also identify some of the essential components for effectively using retrieval, including the importance of within-example in-context learning for acquiring the ability to use information across retrieved examples. In summary, our results illustrate one possible contributor to the relative data inefficiency of current machine learning systems compared to natural intelligence, and help to understand how retrieval methods can complement parametric learning to improve generalization.
@@ 1 Introduction
Figure 1: Conceptual overview of the challenges of using latent information from training experiences, and how retrieval complements parametric learning to overcome them. (1(a)) The reversal curse (Berglund et al., 2024) is an example of how parametric learners, such as language models, consolidate information in ways that depend on the learning task and format. Models that learn a relation in one format can answer queries compatible with the learning format, but not those that reverse the relation—even though the reversed relation is latently implied by the forward one, and the models are fully capable of reversing relations to make inferences in context (Lampinen et al., 2025). (1(b)) Challenges of reversal are one instance of the much broader phenomenon that what is explicitly learned may also latently convey information relevant to other tasks—e.g., multi-hop reasoning, alternative goals, or answering questions in other languages. Like the reversal curse, learning on such sequences may primarily improve performance on the explicit information or goals; however, if the sequence were in context, models would readily be able to make inferences about the latent information. (1(c)) Therefore, explicit retrieval from nonparametric learning systems complements the broader knowledge of parametric learning—by making select, relevant experiences available in context where they can be more flexibly used in ways different from the original task setting in which they were encountered.
@@
When do AI systems fail to generalize like natural intelligence? This question has been central to AI, and has only redoubled with the successes of deep learning over the last decade (e.g., Lake et al., 2017). For example, language models (LMs) show surprising failures to generalize to reversals of their training, e.g., failing to generalize from “X is Y’s son” in training to “Y is X’s parent” at test (Berglund et al., 2024)—though the models can readily make these generalizations in context (Lampinen et al., 2025; cf. (Russin et al., 2025)). Why do models fail to make these generalizations outside the current context, and how could this failure be addressed?
@@
Here, we draw on work in cognitive science to identify a key theme that we believe unifies and generalizes some of these past findings, and points towards a solution. Specifically, we identify the capacity for latent learning (Blodgett, 1929; Tolman, 1948) as a key gap between natural and artificial intelligence. Latent learning is the ability of a system to learn information that is not relevant to the task at hand, but that might be useful for a different future task—which can also be seen as a way of implementing a form of “prospective learning” (De Silva et al., 2023). We suggest that AI systems fail to exhibit latent learning in most cases—that is, AI only learns information insofar as it is relevant to the current task, and only applies information it has learned in the past insofar as it is explicitly or associatively cued by the current task. This is not to say that parametric learning does not generalize well in many cases; indeed, one of the key features of parametric learning is its ability to integrate learning over many experiences to improve generalization within a family of related tasks. However, we will argue that this type of task-driven parametric learning struggles to apply specific information from prior experiences that could be useful in a sufficiently different task at a later time. We will provide several examples and a more formal sketch of this perspective below, unifying a number of previous findings about failures of generalizations and highlighting new ones.
@@
Given this argument, we again take inspiration from cognitive science to suggest a path to a solution. Specifically, in natural intelligence, episodic memory and parametric learning are complementary (McClelland et al., 1995; Kumaran et al., 2016). Episodic memory seems to play an important role in some types of latent learning (Kimble and Greene, 1968; Myers et al., 2000) and the prioritized replay of episodic memories during offline learning separates the parametric knowledge representation from the statistics of the external environment (Kumaran et al., 2016). Moreover, in some of the cases where AI fails to generalize from its parametric learning, it generalizes well in context (Lampinen et al., 2025), suggesting that a system that uses episodic memory to reinstate the relevant information into context could generalize better. Motivated by this intuition, we test a model endowed with an oracle episodic retrieval system during both training and testing, and show that it improves on many of the latent learning failures of current methods. While at some level, this point is just a restatement of the fact that Retrieval Augmented Generation (RAG; Lewis et al., 2020) can be useful, we believe that seeing it in the context of latent learning provides a new perspective on why retrieval is useful, and how current AI systems differ from natural intelligence. It therefore points to an important direction of research for AI: building more effective episodic memory and retrieval systems that approximate the role of the hippocampus as a memory index in natural intelligence (Teyler and DiScenna, 1986).
@@
These results both help to understand some differences between natural and artificial intelligence, suggest new perspectives on extant practices like RAG, and may inspire new research in both cognitive science and AI.
@@ 2.1 Natural intelligence
Latent Learning:  The concept of “latent learning” was introduced by Blodgett (1929), and subsequently elaborated by Tolman (1948), to describe particular phenomena exhibited by rats exploring mazes: the rats seemed to latently learn information that was not relevant to their current goals, but that might be useful in the future. For example, if rats were neither hungry nor thirsty, they would ignore any water or food encountered in the maze. However, if they were later placed in the maze when they were hungry or thirsty, they would rush to the appropriate locations. Thus, even though the rats had not been motivated to seek either food or water when initially exploring the maze, they had latently learned the locations of those resources in a way that allowed them to efficiently exploit them in the future.
@@
Latent learning has been suggested to depend in some cases on the medial temporal lobe (MTL; including the hippocampus) and lesions to this region impair latent learning in some cases (Kimble and Greene, 1968; Myers et al., 2003), as well as other types of generalization (e.g., Bayley and Squire, 2002). (However, note that some latent learning processes may not be MTL-dependent; Kimble et al., 1982; Myers et al., 2000.) More recent computational work has highlighted that hippocampal “preplay” might help to cache the solutions to potential future tasks (Carvalho et al., 2025). Overall, these results suggest that episodic memory may contribute to some—if not all—forms of latent learning in natural intelligence.
@@
Complementary learning systems:  One strand of hippocampal research focuses on “complementary learning systems” (McClelland et al., 1995; Kumaran et al., 2016)—the idea that the rapid learning and episodic memory of the hippocampus provides benefits that complement the slower, more generalized learning within the neocortex. The original argument for why the rapid-learning system is useful was that it could rapidly acquire new information without interfering with broader knowledge in the slower system; salient pieces of rapidly-acquired information could then be consolidated into the slower-learning system over time via replay.
@@
Our work suggests a potential complementary benefit of episodic memory: it could support more flexible use of past experiences in contrast to cortical learning, which might be more closely tied to some aspects of the task or format in which the original experiences occurred. This suggestion is consistent with the lesion findings above, and more broadly compatible with arguments that the hippocampus plays a crucial role in organizing (spatial) knowledge into structures that support flexible learning and behavior (Behrens et al., 2018; Raju et al., 2024).
@@
Transformers, context, and episodic memory in natural intelligence:  Several recent works note related connections between key computational features of transformers and natural intelligence. Most saliently, Russin et al. (2025) highlight how capacity and flexibility tradeoffs between parametric learning and in-context learning relate to similar tradeoffs in human intelligence. More generally, various works have made computational connections between transformer attention mechanisms and episodic memory in the brain (Whittington et al., 2022; Gershman et al., 2025). Our work complements these, by showing how these connections could point to important contributions of episodic memory to generalization, in both transformers and natural intelligence.
@@ 2.2 Artificial Intelligence
Prospective learning:  The topic of latent learning also relates to the argument of De Silva et al. (2023) that machine learning assumes that the future will be identical in distribution to the past, but natural intelligence learns more prospectively, in a way that is oriented towards generalizing in the future. The challenges of latent learning can be interpreted as a type of prospective learning—learning information not because it is useful under the current task distribution, but because it might be useful if the task distribution changes in the future. However, the type of prospective learning focused on by De Silva et al. involves explicitly modeling how tasks might change in the future; retrieval avoids this challenge by leaving the question of which tasks may be relevant to the future (see below).
Meta-learning and in-context learning in transformers:  Another way a system can adapt to a novel future task is by learning that task once it appears; many works have explored how this capability can be meta-learned (e.g., Santoro et al., 2016; Wang et al., 2017), often using transformers as a memory in context (Parisotto et al., 2020; Lampinen et al., 2021). Many works have similarly interpreted the emergent in-context learning ability in language models (Brown et al., 2020) as a type of implicit meta-learning over the training distribution (Xie et al., 2022; Chan et al., 2022a; Lampinen et al., 2024). Thus, there has been substantial interest in meta-learning (across contexts) how to learn within a single context. However, meta-learning alone does not fully address how knowledge from past contexts could be directly applied at present.
Using memory of prior contexts in agents & language models:  Several prior works in RL focused on the benefit of retrieving memory of past contexts to help adapt to present ones (e.g., Ritter et al., 2018; Pritzel et al., 2017; Goyal et al., 2022). More recently, there has been substantial interest in retrieving related training documents to improve test time performance in LMs. Seminal work on Retrieval Augmented Generation (RAG; Lewis et al., 2020) showed that retrieving Wikipedia articles, and adding them to the context [^1] The full RAG strategy involves additional marginalization over multiple documents, etc.; however, each document is effectively just prepended to the context., could substantially improve language model responses. More recent works have used retrieved documents to fine-tune the models (Hardt and Sun, 2024; Hübotter et al., 2025) or explored incorporating retrieval into the context throughout training (e.g., Khattab and Zaharia, 2020; Borgeaud et al., 2022).
@@
Patterns of generalization in language models:  Several streams of work identifying failures and successes of generalization in LMs helped inspire our investigations. First, the Reversal Curse (Berglund et al., 2024), and failures of multi-hop inference (Balesni et al., 2024), are intriguing failures of LMs to generalize fully the information available in their fine-tuning data. Other studies similarly find inadequate cross-linguistic transfer of knowledge in language models (Qi et al., 2023; Aggarwal et al., 2025; Goldman et al., 2025). We interpret these findings as instances of the broader phenomenon we highlight: language models learn what is explicit in the training data, in ways that may not generalize to all its latent implications for other questions (see Fig. 1(b)).
However, other works have found instances of surprisingly-successful generalization in language models. For example, several works demonstrate cases where models seem to generalize to a test example based on facts stated in other training documents (Berglund et al., 2023; Meinke and Evans, 2023). Similarly, Cook et al. (2025) show that models can generalize from code for functions in training to execute those functions at test time. These examples show cases where language models do exhibit surprising generalization “out-of-context” in a way reminiscent of latent learning, though the absolute generalization performance observed is usually low. We return to this issue in the discussion.
Generalization from fine-tuning and in-context learning:  Several closely-related works show differences in generalization between in-context learning and parametric learning in pretrained language models (Lampinen et al., 2025; Park et al., 2025; cf. Chan et al., 2022b). Like our results, these works show that context-based learning can offer unique contributions beyond parametric learning. However, they focus on pretrained language models and narrower evaluations, and do not clearly identify the division between explicit and latent knowledge that we highlight here. Many of these works propose data-augmentation-based solution, as do other related approaches (Akyürek et al., 2024; Yang et al., 2025; Ruan et al., 2025). Augmentation can enable expending additional train-time inferences to improve test-time generalization by filling in the latent information; however, this requires the model (or model trainer) to infer a priori which latent tasks should be targeted for augmentation. Our results show an alternative perspective by which retrieval from episodic memory can allow performance without needing to decide at training time which tasks may be relevant later—overcoming similar challenges while remaining more flexible. We return to these issues in the discussion.
@@ 4.2 Simple reversals
This dataset is loosely inspired by the reversal curse dataset of (Berglund et al., 2024), and the simple reversals dataset of Lampinen et al. (2025), but is designed to have sufficient diversity for (minimal) training from scratch. To do so, we generated a set of facts about 1,000 entities and 20 relations (plus a reverse relation for each, for a total of 40). e.g., if X and Y are entities, a relation might be “X contains Y” and its reverse might be “Y subset_of X.” We train the model on a set of 20,000 relations in the forward direction, and 19,800 of their reversals, holding out 200 reverse relations for a test set. To create training documents, we augment each relation sentence with up to 5 random prefix and suffix tokens. We also include 12.5% of ICL sequences that have both forward and backward relations in the context. We validate generalization to the trained forward relations with novel prefixes, and test generalization to the held-out reversals.

### v2

@@ Abstract
When do machine learning systems fail to generalize, and what mechanisms could improve their generalization? Here, we draw inspiration from cognitive science to argue that one weakness of parametric machine learning systems is their failure to exhibit latent learning—learning information that is not relevant to the task at hand, but that might be useful in a future task. We show how this perspective links failures ranging from the reversal curse in language modeling to new findings on agent-based navigation. We then highlight how cognitive science points to episodic memory as a potential part of the solution to these issues. Correspondingly, we show that a system with an oracle retrieval mechanism can use learning experiences more flexibly to generalize better across many of these challenges. We also identify some of the essential components for effectively using retrieval, including the importance of within-example in-context learning for acquiring the ability to use information across retrieved examples. In summary, our results illustrate one possible contributor to the relative data inefficiency of current machine learning systems compared to natural intelligence, and help to understand how retrieval methods can complement parametric learning to improve generalization. We close by discussing some of the links between these findings and prior results in cognitive science and neuroscience, and the broader implications.
@@ 1 Introduction
Figure 1: Conceptual overview of the challenges of using latent information from training experiences, and how retrieval complements parametric learning to overcome them. (1(a)) The reversal curse (berglund2024reversal) is an example of how parametric learners, such as language models, consolidate information in ways that depend on the learning task and format. Models that learn a relation in one format can answer queries compatible with the learning format, but not those that reverse the relation—even though the reversed relation is latently implied by the forward one, and the models are fully capable of reversing relations to make inferences in context (lampinen2025generalization). (1(b)) Challenges of reversal are one instance of the much broader phenomenon that what is explicitly learned may also latently convey information relevant to other tasks—e.g., multi-hop reasoning, alternative goals, or answering questions in other languages. Like the reversal curse, learning on such sequences may primarily improve performance on the explicit information or goals; however, if the sequence were in context, models would readily be able to make inferences about the latent information. (1(c)) Therefore, explicit retrieval of specific experiences from nonparametric learning systems complements the broader knowledge of parametric learning—by making select, relevant experiences available in context where the latent information they contain can be used more flexibly, in ways different from the original task setting in which they were encountered. (1(d)) Thus, we will typically expect systems with solely parametric learning to perform well at new tests of knowledge that is explicit in learning experiences, but we expect selective performance advantages for episodic retrieval in tests of knowledge or tasks that are latent in learning experiences—as we demonstrate below. (These illustrative results are adapted from the simple reversals experiments below.)
@@
When do AI systems fail to generalize like natural intelligence? This question has been central to AI, and has only redoubled with the successes of deep learning over the last decade (e.g., lake2017building). For example, language models (LMs) show surprising failures to generalize to reversals of their training, e.g., failing to generalize from “X is Y’s son” in training to “Y is X’s parent” at test (berglund2024reversal)—though the models can readily make these generalizations in context (lampinen2025generalization; cf. russin2025parallel). Why do models fail to make these generalizations outside the current context, and how could this failure be addressed?
@@
Here, we draw on work in cognitive science to identify a key theme that we believe unifies and generalizes some of these past findings, and points towards a solution. Specifically, we identify the capacity for latent learning (blodgett1929effect; tolman1948cognitive) as a key gap between natural and artificial intelligence. Latent learning is the ability of a system to learn information that is not relevant to the task at hand, but that might be useful for a different future task—which can also be seen as a way of implementing a form of “prospective learning” (desilva2023prospective). We suggest that AI systems fail to exhibit latent learning in most cases—that is, AI only learns information insofar as it is relevant to the current task, and only applies information it has learned in the past insofar as it is explicitly or associatively cued by the current task. This is not to say that parametric learning does not generalize well in many cases; indeed, one of the key features of parametric learning is its ability to integrate learning over many experiences to improve generalization within a family of related tasks. However, we will argue that this type of task-driven parametric learning struggles to apply specific information from prior experiences that could be useful in a sufficiently different task at a later time. We will provide several examples and a more formal sketch of this perspective below, unifying a number of previous findings about failures of generalization and highlighting new ones.
@@
Given this argument, we again take inspiration from cognitive science to suggest a path to a solution. Specifically, in natural intelligence, episodic memory and parametric learning are complementary (mcclelland1995there; kumaran2016learning). Episodic memory seems to play an important role in some types of latent learning (kimble1968absence; myers2000latent) and other types of generalization (bayley2002medial; eichenbaum2009neurobiology). Moreover, in some of the cases where AI fails to generalize from its parametric learning, it generalizes well in context (lampinen2025generalization), suggesting that a system that uses episodic memory to reinstate the relevant information into context could generalize better. Motivated by this intuition, we test a model endowed with an oracle episodic retrieval system during both training and testing, and show that it improves on many of the latent learning failures of current methods. While at some level, this point is just a restatement of the fact that Retrieval Augmented Generation (RAG; lewis2020retrieval) can be useful, we believe that seeing it in the context of latent learning provides a new perspective on why retrieval is useful, and how current AI systems differ from natural intelligence. It therefore points to an important direction of research for AI: building more effective episodic memory and retrieval systems that approximate the role of the hippocampus in enabling generalization.
@@
These results shed light on some differences between natural and artificial intelligence, suggest new perspectives on extant practices like RAG, and may inspire new research in both cognitive science and AI.
@@ 2.1 Natural intelligence
Latent Learning:  The concept of “latent learning” was introduced by blodgett1929effect, and subsequently elaborated by tolman1948cognitive, to describe particular phenomena exhibited by rats exploring mazes: the rats seemed to latently learn information that was not relevant to their current goals, but that might be useful in the future. For example, if rats were neither hungry nor thirsty, they would ignore any water or food encountered in the maze. However, if they were later placed in the maze when they were hungry or thirsty, they would rush to the appropriate locations. Thus, even though the rats had not been motivated to seek either food or water when initially exploring the maze, they had latently learned the locations of those resources in a way that allowed them to efficiently exploit them in the future.
@@
Latent learning has been suggested to depend in some cases on the medial temporal lobe (MTL; including the hippocampus) and lesions to this region impair latent learning in some cases (kimble1968absence; myers2003dissociating), as well as other types of generalization (e.g., bayley2002medial; eichenbaum2009neurobiology). (However, note that some latent learning processes may not be MTL-dependent; kimble1982further; myers2000latent.) More recent computational work has highlighted that hippocampal “preplay” might help to cache the solutions to potential future tasks (carvalho2025preemptive). Overall, these results suggest that episodic memory may contribute to some—if not all—forms of latent learning in natural intelligence.
@@
Complementary learning systems:  One strand of hippocampal research focuses on “complementary learning systems” (mcclelland1995there; kumaran2016learning)—the idea that the rapid learning and episodic memory of the hippocampus provides benefits that complement the slower, more generalized learning within the neocortex. The original argument for why the rapid-learning system is useful was that it could rapidly acquire new information without interfering with broader knowledge in the slower system; salient pieces of rapidly-acquired information could then be consolidated into the slower-learning system over time via replay.
@@
Our work suggests a potential complementary benefit of episodic memory: it could support more flexible use of past experiences in contrast to cortical learning, while cortical learning might be more closely tied to some aspects of the task or format in which the original experiences occurred. This suggestion is consistent with the lesion findings above, and more broadly compatible with arguments that the hippocampus plays a crucial role in organizing (spatial) knowledge into structures that support flexible learning and behavior (behrens2018cognitive; raju2024space)—and in linking memories together to enable generalizations such as transitive inferences (eichenbaum2009neurobiology; kumaran2012generalization) and imagination (hassabis2007patients).
@@
Transformers, context, and episodic memory in natural intelligence:  Several recent works note related connections between key computational features of transformers and natural intelligence. Most saliently, russin2025parallel highlight how capacity and flexibility tradeoffs between parametric learning and in-context learning relate to similar tradeoffs in human intelligence. More generally, various works have made computational connections between transformer attention mechanisms and episodic memory in the brain (whittington2022relating; gershman2025key). Our work complements these, by showing how these connections could point to important contributions of episodic memory to generalization, in both transformers and natural intelligence.
@@ 2.2 Artificial Intelligence
Prospective learning:  The topic of latent learning also relates to the argument of desilva2023prospective that machine learning assumes that the future will be identical in distribution to the past, but natural intelligence learns more prospectively, in a way that is oriented towards generalizing in the future. The challenges of latent learning can be interpreted as a type of prospective learning—learning information not because it is useful under the current task distribution, but because it might be useful if the task distribution changes in the future. However, the type of prospective learning focused on by De Silva et al. involves explicitly modeling how tasks might change in the future; retrieval avoids this challenge by leaving the question of which tasks may be relevant to the future (see below).
Meta-learning and in-context learning in transformers:  Another way a system can adapt to a novel future task is by learning that task once it appears; many works have explored how this capability can be meta-learned (e.g., santoro2016meta; wang2017learning), often using transformers as a memory in context (parisotto2020stabilizing; lampinen2021towards). Many works have similarly interpreted the emergent in-context learning ability in language models (brown2020language) as a type of implicit meta-learning over the training distribution (xie2022an; chan2022data; lampinen2024broader). Thus, there has been substantial interest in meta-learning (across contexts) how to learn within a single context. However, meta-learning alone does not fully address how knowledge from past contexts could be directly applied at present.
Using memory of prior contexts in agents & language models:  Several prior works in RL focused on the benefit of retrieving memory of past contexts to help adapt to present ones (e.g., ritter18been; pritzel2017neural; goyal2022retrieval; humphreys2022large). More recently, there has been substantial interest in retrieving related training documents to improve test time performance in LMs. Seminal work on Retrieval Augmented Generation (RAG; lewis2020retrieval) showed that retrieving Wikipedia articles, and adding them to the context [^1] The full RAG strategy involves additional marginalization over multiple documents, etc.; however, each document is effectively just prepended to the context., could substantially improve language model responses. More recent works have used retrieved documents to fine-tune the models (hardt2024testtime; hubotter2025efficiently) or explored incorporating retrieval into the context throughout training (e.g., khattab2020colbert; borgeaud2022improving).
@@
Patterns of generalization in language models:  Several streams of work identifying failures and successes of generalization in LMs helped inspire our investigations. First, the Reversal Curse (berglund2024reversal), and failures of multi-hop inference (balesni2024two), are intriguing failures of LMs to generalize fully the information available in their fine-tuning data. Other studies similarly find inadequate cross-linguistic transfer of knowledge in language models (qi2023cross; aggarwal2025language; goldman2025eclektic). We interpret these findings as instances of the broader phenomenon we highlight: language models learn what is explicit in the training data, in ways that may not generalize to all its latent implications for other questions (see Fig. 1(b)).
However, other works have found instances of surprisingly-successful generalization in language models. For example, several works demonstrate cases where models seem to generalize to a test example based on facts stated in other training documents (berglund2023taken; meinke2023tell). Similarly, cook2025programming show that models can generalize from code for functions in training to execute those functions at test time. These examples show cases where language models do exhibit surprising generalization “out-of-context” in a way reminiscent of latent learning, though the absolute generalization performance observed is usually low. We return to this issue in the discussion.
Generalization from fine-tuning and in-context learning:  Several closely-related works show differences in generalization between in-context learning and parametric learning in pretrained language models (lampinen2025generalization; park2025new; cf. chan2022transformers). Like our results, these works show that context-based learning can offer unique contributions beyond parametric learning. However, they focus on pretrained language models and narrower evaluations, and do not clearly identify the division between explicit and latent knowledge that we highlight here. Many of these works propose data-augmentation-based solution, as do other related approaches (akyurek2024deductive; yang2025synthetic; ruan2025reasoning). Augmentation can enable expending additional train-time inferences to improve test-time generalization by filling in the latent information; however, this requires the model (or model trainer) to infer a priori which latent tasks should be targeted for augmentation. Our results show an alternative perspective by which retrieval from episodic memory can allow performance without needing to decide at training time which tasks may be relevant later—overcoming similar challenges while remaining more flexible. We return to these issues in the discussion.
@@ 4.2 Simple reversals
This dataset is loosely inspired by the reversal curse dataset of (berglund2024reversal), and the simple reversals dataset of lampinen2025generalization, but is designed to have sufficient diversity for (minimal) training from scratch. To do so, we generated a set of facts about 1,000 entities and 20 relations (plus a reverse relation for each, for a total of 40). e.g., if X and Y are entities, a relation might be “X contains Y” and its reverse might be “Y subset_of X.” We train the model on a set of 20,000 relations in the forward direction, and 19,800 of their reversals, holding out 200 reverse relations for a test set. To create training documents, we augment each relation sentence with up to 5 random prefix and suffix tokens. We also include 12.5% of ICL sequences that have both forward and backward relations in the context. We validate generalization to the trained forward relations with novel prefixes, and test generalization to the held-out reversals.